leaqueen:

The dot plots below show the ages of students belonging to two groups of salsa classes: A dot plot shows two divisions labeled Group A and Group B. The horizontal axis is labeled as Age of Salsa Students in years. Group A shows 3 dots at 5, 4 dots at 10, 6 dots at 17, 4 dots at 24, and 3 dots at 28. Group B shows 6 dots at 7, 3 dots at 10, 4 dots at 14, 5 dots at 17, and 2 dots at 22. Based on visual inspection, which group most likely has a lower mean age of salsa students? Explain your answer using two or three sentences. Make sure to use facts to support your answer.

YSNJayy:

Based on visual inspection, Group B most likely has a lower mean age of salsa students. As we can see from the dot plot, the age of Group A is evenly distributed at different ages, however, Group B has a tendency of having more students at lower age below 17. While Group A has 13 students who are above or at the age of 17, only 2 students in Group B are above 17. Although in Group A there are very young students, the students who are at very old age are likely to pull up the mean age. Therefore, group B is more likely to have a lower mean age. To check if it is correct: Mean age of Group A =(3x5 + 4x10 + 6x17 + 4x24 + 3x28)÷(3+4+6+4+3) =337÷20 =16.85 Mean age of group B = (6x7 + 3x10 + 4x14 + 5x17 + 2x22)÷(6+3+4+5+2) =257÷20 =12.85 Therefore, 12.85(Group A) < 16.85(Group B), the answer is Group B.

leaqueen:

thx your the best

YSNJayy:

np